# Decentralized Smoothing ADMM for Quantile Regression with Non-Convex Sparse Penalties

### v1

@@ Abstract
In the rapidly evolving internet-of-things (IoT) ecosystem, effective data analysis techniques are crucial for handling distributed data generated by sensors. Addressing the limitations of existing methods, such as the sub-gradient approach, which fails to distinguish between active and non-active coefficients effectively, this paper introduces the decentralized smoothing alternating direction method of multipliers (DSAD) for penalized quantile regression. Our method leverages non-convex sparse penalties like the minimax concave penalty (MCP) and smoothly clipped absolute deviation (SCAD), improving the identification and retention of significant predictors. DSAD incorporates a total variation norm within a smoothing ADMM framework, achieving consensus among distributed nodes and ensuring uniform model performance across disparate data sources. This approach overcomes traditional convergence challenges associated with non-convex penalties in decentralized settings. We present theoretical proofs and extensive simulation results to validate the effectiveness of the DSAD, demonstrating its superiority in achieving reliable convergence and enhancing estimation accuracy compared with prior methods.
@@ I Introduction
  • The demonstration of DSAD’s capability to effectively manage the sparsity and accuracy challenges posed by non-convex penalties, achieving node consensus and robust learning in a decentralized setting.
  • Empirical validation of DSAD through simulations, showcasing superior performance and setting a new benchmark in distributed, penalized quantile regression.
@@ II Problem formulation
Consider a scalar response variable $Y$ and a $P$-dimensional vector of predictors $\mathbf{x}$. The conditional cumulative distribution function is defined as $F_{Y}(y|\mathbf{x})=P(Y\leq y|\mathbf{x})$, and the $\tau$-th conditional quantile, for $\tau\in(0,1)$, is given by $Q_{Y}(\tau|\mathbf{x})=\inf\{y:F_{Y}(y|\mathbf{x})\geq\tau\}$. In quantile regression, this quantile can be linearly modeled as:
$Q_{Y}(\tau|\mathbf{x})=\mathbf{x}^{\text{T}}\boldsymbol{\beta}_{\tau}+q_{\tau}^{\epsilon},$ (1)
where $\boldsymbol{\beta}_{\tau}$ are the regression coefficients specific to quantile $\tau$, and $q_{\tau}^{\epsilon}$ represents the $\tau$-th quantile of the noise [15].
@@
$\hat{\mathbf{w}}=\operatorname*{arg\,min}_{\mathbf{w}}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-\mathbf{\bar{x}}_{i}^{\text{T}}\mathbf{w})+P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}\right),$ (2)
@@
$\min_{\{\mathbf{w}_{l},\mathbf{z}_{l}\}_{l=1}^{L}}$ $\sum_{l=1}^{L}\left(\frac{1}{2}\left(\|\mathbf{z}_{l}\|_{1}+(2\tau-1)\mathbf{1}_{n}^{\text{T}}\mathbf{z}_{l}\right)+nP_{\lambda,\gamma}(\mathbf{w}_{l})\right),$
subject to $\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}=\mathbf{y}^{(l)},\forall l\in\{1,\cdots,L\},$ (3)
$\mathbf{w}_{l}=\mathbf{w}_{j},\forall j\in\mathcal{N}_{l}.$ (4)
@@
$\min_{\{\mathbf{w}_{l},\mathbf{z}_{l}\}_{l=1}^{L}}$ $\sum_{l=1}^{L}\frac{1}{2}\mathopen{}\left(\|\mathbf{z}_{l}\|_{1}+\mathopen{}\left(2\tau-1\right)\mathclose{}\mathbf{1}_{n}^{\text{T}}\mathbf{z}_{l}\right)\mathclose{},$
@@
$\mathbf{g}_{lj}=\mathbf{w}_{l},\quad\mathbf{g}_{jl}=\mathbf{w}_{j},\forall j\in\mathcal{N}_{l},j>l.$ (5)
@@
III Smoothing Decentraizled ADMM
To enhance ADMM framework optimization of non-smooth functions, smoothing techniques transform $g$ into smooth functions $\tilde{g}$, suitable for locally Lipschitz environments, improving optimization manageability through properties like continuity, differentiability, and beneficial gradient properties, ensuring $\tilde{g}(\mathbf{x},\mu)$ converges to $g(\mathbf{x})$ as $\mu\to 0^{+}$ [22].
Our smoothing ADMM algorithm employs these properties to approximate the $\|\mathbf{z}\|_{1}$ norm effectively $\|\mathbf{z}\|_{1}$ norm by using a sum of smooth functions for each $|z_{i}|$ presented in [22] as $h\mathopen{}\left(\mathbf{z},\mu\right)\mathclose{}=\sum_{i=1}^{n}f\mathopen{}\left(z_{i},\mu\right)\mathclose{}$, where:
@@ III Smoothing Decentraizled ADMM
¯${\bar{\mathcal{L}}}_{\sigma_{\Psi},\sigma_{\xi},\mu}\mathopen{}\left(\mathbf{W},\mathbf{Z},\mathbf{G},\boldsymbol{\Psi},\boldsymbol{\xi}\right)\mathclose{}=\sum_{l=1}^{L}\Bigg{(}\frac{1}{2}\mathopen{}h\left(\mathbf{z}_{l},\mu\right)+\tau\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}-\\ \frac{1}{2}\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}+M_{l}P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}_{l}\right)\mathclose{}+\frac{\sigma_{\Psi}}{2}\mathopen{}\left\|\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}-\mathbf{y}^{(l)}+\frac{\boldsymbol{\Psi}_{l}}{\sigma_{\Psi}}\right\|_{2}^{2}\mathclose{}+\\ \sum_{j\in\mathcal{N}_{l},j>l}\bigg{(}\omega\hskip 1.42262pth\left(\mathbf{g}_{lj}-\mathbf{g}_{jl},\mu\right)+\frac{\sigma_{\xi}}{2}\|\mathbf{w}_{i}-\mathbf{g}_{lj}\|^{2}_{2}+\frac{\sigma_{\xi}}{2}\|\mathbf{w}_{j}-\mathbf{g}_{jl}\|^{2}_{2}\\ +\boldsymbol{\xi}_{lj}^{\text{T}}(\mathbf{w}_{l}-\mathbf{g}_{lj})+\boldsymbol{\xi}_{jl}^{\text{T}}(\mathbf{w}_{j}-\mathbf{g}_{jl})\bigg{)}\Bigg{)},$ (7)
@@
To iteratively purify the approximation, we update the parameters $\mu$, $\sigma_{\Psi}$, $\sigma_{\xi}$ in each iteration using the following rules, where $c>0$, $d>0$ and $\beta>0$:
@@
The update process for each $\mathbf{w}_{l}$ is divided into $P+1$ sequential steps, with the $p$-th element updated in its corresponding step. After simplifications in each agent $l$, the update formula for the $p$-th for $p\leq P$ is as follows:
$w_{p,l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\textbf{Prox}_{g_{\lambda,\gamma}}\mathopen{}\left(\frac{a_{p,l}}{\Upsilon^{(l)}_{p}};\frac{M_{l}}{\Upsilon^{(l)}_{p}}\right)\mathclose{},$ (9)
@@
$a_{p,l}=-\left(\mathbf{x}^{(l)}_{:,p}\right)^{\text{T}}\left(\mathbf{X}^{(l)}_{:,<p}\mathbf{w}^{\mathopen{}\left(k+1\right)\mathclose{}}_{<p,l}+\mathbf{X}^{(l)}_{:,>p}\mathbf{w}^{\mathopen{}\left(k\right)\mathclose{}}_{>p,l}\right)\\ -\left(\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)^{\text{T}}\mathbf{X}^{(l)}_{:,p}+\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\left(\mathbf{y}^{(l)}-\mathbf{z}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)^{\text{T}}\mathbf{X}^{(l)}_{:,p}+\\ \sum_{j\in\mathcal{N}_{l},j<l}(\sigma_{\xi}^{(k+1)}g_{lj,p}^{(k)}-\xi_{lj,p}^{(k)})+\sum_{j\in\mathcal{N}_{l},j>l}(\sigma_{\xi}^{(k+1)}g_{jl,p}^{(k)}-\xi_{jl,p}^{(k)}),$ (10)
@@
$\Upsilon^{(l)}_{p}=\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\mathopen{}\left\|\mathbf{X}^{(l)}_{:,p}\right\|_{2}^{2}\mathclose{}+\kappa_{\xi}^{(k+1)}.$ (11)
Both MCP and SCAD admit closed-form solutions of the proximal operator [23]. For $P+1$-th element as the penalty function is zero we have:
$w_{P+1,l}=\frac{a_{P+1,l}}{\Upsilon^{(l)}_{P+1}}.$ (12)
@@
$\mathbf{z}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\operatorname*{arg\,min}_{\mathbf{z}_{l}}\frac{1}{2}\mathopen{}\left(\sum_{j=1}^{n}f\mathopen{}\left(z_{l,j},\mu^{\mathopen{}\left(k+1\right)\mathclose{}}\right)\mathclose{}+\mathopen{}\left(2\tau-1\right)\mathclose{}\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}\right)\mathclose{}$
$+\mathopen{}\left(\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)\mathclose{}^{\text{T}}\mathbf{z}_{l}+\frac{\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}}{2}\mathopen{}\left\|\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}-\mathbf{y}^{(l)}\right\|_{2}^{2}\mathclose{}.$ (13)
It can be shown that the update step of each $\mathbf{z}_{l}$ in ADMM has a closed-form solution. By merging $\mathopen{}\left(\tau-\frac{1}{2}\right)\mathclose{}\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}$, $\mathopen{}\left(\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)\mathclose{}^{\text{T}}\mathbf{z}_{l}$, and $\mathopen{}\left\|\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}-\mathbf{y}^{(l)}\right\|_{2}^{2}\mathclose{}$ together, a component-wise solution can be obtained as
$\mathbf{z}_{i}^{\mathopen{}\left(k+1\right)\mathclose{}}=\textbf{Prox}_{f\mathopen{}\left(\cdot,\mu^{\mathopen{}\left(k+1\right)\mathclose{}}\right)\mathclose{}}\mathopen{}\left(\alpha_{l,i};\frac{\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}}{2}\right)_{i=1}^{M_{l}}\mathclose{},$ (14)
@@
Updates for both $\mathbf{g}_{lj}$ and $\mathbf{g}_{jl}$ for $j\in\mathcal{N}_{l}$, $j<l$ can be performed concurrently as:
$\begin{bmatrix}\mathbf{g}_{lj}^{(k+1)}\\ \mathbf{g}_{jl}^{(k+1)}\end{bmatrix}=\operatorname*{arg\,min}_{\mathbf{g}_{lj},\mathbf{g}_{jl}}\omega\sum_{p=1}^{P+1}f({g}_{lj,p}-{g}_{jl,p},\mu^{\mathopen{}\left(k+1\right)\mathclose{}})+\frac{\sigma_{\xi}}{2}\Bigg{(}\\ \left\|\mathbf{g}_{lj}-\mathbf{w}_{l}^{(k+1)}+\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}^{(k+1)}}\right\|^{2}_{2}+\left\|\mathbf{g}_{jl}-\mathbf{w}_{l}^{(k+1)}+\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}}\right\|^{2}_{2}\Bigg{)}.$ (16)
@@
$\begin{bmatrix}\mathbf{g}_{lj}^{(k+1)}\\ \mathbf{g}_{jl}^{(k+1)}\end{bmatrix}=\\ \frac{1}{2}\begin{bmatrix}\mathbf{w}_{l}^{(k+1)}-\frac{\boldsymbol{\xi}_{lj}^{(k)}}{\sigma_{\xi}^{(k+1)}}+\mathbf{w}_{j}^{(k+1)}-\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}^{(k+1)}}\\ \mathbf{w}_{i}^{(k+1)}-\frac{\boldsymbol{\xi}_{lj}^{(k)}}{\sigma_{\xi}^{(k+1)}}+\mathbf{w}_{j}^{(k+1)}-\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\kappa_{\xi}^{(k+1)}}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-\mathbf{e}\\ \mathbf{e}\end{bmatrix},$ (17)
@@
$\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}+\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\mathopen{}\left(\mathbf{z}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}+\mathbf{X}^{(l)}\mathbf{w}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}-\mathbf{y}^{(l)}\right)\mathclose{},$ (18)
$\boldsymbol{\xi}^{(k+1)}_{lj}=\boldsymbol{\xi}^{(k)}_{lj}+\sigma^{(k+1)}_{\xi}\left(\mathbf{w}^{(k+1)}_{l}-\mathbf{g}^{(k+1)}_{lj}\right),$ (19)
$\boldsymbol{\xi}^{(k+1)}_{jl}=\boldsymbol{\xi}^{(k)}_{jl}+\sigma_{\xi}^{(k+1)}\left(\mathbf{w}^{(k+1)}_{j}-\mathbf{g}^{(k+1)}_{jl}\right).$ (20)
The proposed ADMM-based method for solving the decentralized sparse-penalized quantile regression is summarized in Algorithm 1 and the convergence is provided in Theorem 1.
@@
$\mathbf{X}^{\textbf{T}}\boldsymbol{\Psi}^{*}\in n\partial P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}^{*}\right)\mathclose{},$ (21a)
$\boldsymbol{\Psi}^{*}\in\partial\sum_{i=1}^{n}\rho_{\tau}\mathopen{}\left(z_{i}^{*}\right)\mathclose{},$ (21b)
$\mathbf{z}^{*}+\mathbf{X}\mathbf{w}^{*}-\mathbf{y}=0.$ (21c)
@@
¯${\bar{\mathcal{L}}}_{\sigma_{\Psi}^{(K^{\prime})},\sigma_{\xi}^{(K^{\prime})},\mu^{(K^{\prime})}}\mathopen{}\left(\mathbf{W}^{(K^{\prime})},\mathbf{Z}^{(K^{\prime})},\mathbf{G}^{(K^{\prime})},\boldsymbol{\Psi}^{(K^{\prime})},\boldsymbol{\xi}^{(K^{\prime})}\right)\mathclose{}-\\ {\bar{\mathcal{L}}}_{\sigma_{\Psi}^{(K)},\sigma_{\xi}^{(K)},\mu^{(K)}}\mathopen{}\left(\mathbf{W}^{(K)},\mathbf{Z}^{(K)},\mathbf{G}^{(K)},\boldsymbol{\Psi}^{(K)},\boldsymbol{\xi}^{(K)}\right)\mathclose{}\leq\\ -S_{K^{\prime}}-S^{\prime}_{K^{\prime}}+D,$ (22)
@@
$S_{K^{\prime}}=\sum_{k=K}^{K^{\prime}-1}\sum_{l=1}^{L}\eta^{(k+1)}\|\mathbf{z}_{l}^{(k+1)}-\mathbf{z}_{l}^{(k)}\|_{2}^{2}\\ +\varkappa^{(k+1)}\|\mathbf{w}_{l}^{(k+1)}-\mathbf{w}_{l}^{(k)}\|_{2}^{2},$ (23)
@@
$S_{K^{\prime}}=\sum_{k=K}^{K^{\prime}-1}\sum_{l=1}^{L}\sum_{j\in\mathcal{N}_{l},j>l}\iota^{(k+1)}\|\mathbf{q}_{lj}^{(k+1)}-\mathbf{q}_{lj}^{(k)}\|^{2}_{2},$ (24)

### v2

@@ Abstract
In the rapidly evolving internet-of-things (IoT) ecosystem, effective data analysis techniques are crucial for handling distributed data generated by sensors. Addressing the limitations of existing methods, such as the sub-gradient approach, which fails to distinguish between active and non-active coefficients effectively, this paper introduces the decentralized smoothing alternating direction method of multipliers (DSAD) for penalized quantile regression. Our method leverages non-convex sparse penalties like the minimax concave penalty (MCP) and smoothly clipped absolute deviation (SCAD), improving the identification and retention of significant predictors. DSAD incorporates a total variation norm within a smoothing ADMM framework, achieving consensus among distributed nodes and ensuring uniform model performance across disparate data sources. This approach overcomes traditional convergence challenges associated with non-convex penalties in decentralized settings. We present convergence proof and extensive simulation results to validate the effectiveness of the DSAD, demonstrating its superiority in achieving reliable convergence and enhancing estimation accuracy compared with prior methods.
@@ I Introduction
  • The demonstration of DSAD’s capability to effectively manage the sparsity and accuracy challenges posed by non-convex penalties, achieving node consensus and local optimality in a decentralized setting.
  • Empirical validation of DSAD through simulations, showcasing superior performance and setting a new benchmark in distributed penalized quantile regression.
@@ II Problem formulation
Consider a scalar response variable $Y$ and a $P$-dimensional vector of predictors $\mathbf{x}$. The conditional cumulative distribution function is defined as $F_{Y}(y|\mathbf{x})=P(Y\leq y|\mathbf{x})$, and the $\tau$th conditional quantile, for $\tau\in(0,1)$, is given by $Q_{Y}(\tau|\mathbf{x})=\inf\{y:F_{Y}(y|\mathbf{x})\geq\tau\}$. In quantile regression, this quantile can be linearly modeled as:
$Q_{Y}(\tau|\mathbf{x})=\mathbf{x}^{\text{T}}\boldsymbol{\beta}_{\tau}+q_{\tau}^{\epsilon}$ (1)
where $\boldsymbol{\beta}_{\tau}$ are the regression coefficients specific to quantile $\tau$, and $q_{\tau}^{\epsilon}$ represents the $\tau$th quantile of the noise [15].
@@
$\hat{\mathbf{w}}=\operatorname*{arg\,min}_{\mathbf{w}}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-\mathbf{\bar{x}}_{i}^{\text{T}}\mathbf{w})+P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}\right)$ (2)
@@
$\min_{\{\mathbf{w}_{l},\mathbf{z}_{l}\}_{l=1}^{L}}$ $\sum_{l=1}^{L}\left(\frac{1}{2}\left(\|\mathbf{z}_{l}\|_{1}+(2\tau-1)\mathbf{1}_{n}^{\text{T}}\mathbf{z}_{l}\right)+nP_{\lambda,\gamma}(\mathbf{w}_{l})\right)$
subject to $\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}=\mathbf{y}^{(l)},\forall l\in\{1,\cdots,L\}$ (3)
$\mathbf{w}_{l}=\mathbf{w}_{j}\hskip 5.69054pt\forall j\in\mathcal{N}_{l}$ (4)
@@
$\min_{\{\mathbf{w}_{l},\mathbf{z}_{l}\}_{l=1}^{L}}$ $\sum_{l=1}^{L}\frac{1}{2}\mathopen{}\left(\|\mathbf{z}_{l}\|_{1}+\mathopen{}\left(2\tau-1\right)\mathclose{}\mathbf{1}_{n}^{\text{T}}\mathbf{z}_{l}\right)\mathclose{}$
@@
$\mathbf{g}_{lj}=\mathbf{w}_{l},\quad\mathbf{g}_{jl}=\mathbf{w}_{j},\forall j\in\mathcal{N}_{l},j>l$ (5)
@@
III Decentralized Smoothing ADMM
To enhance ADMM framework optimization of non-smooth functions, smoothing techniques transform a function $g$ into a series of smooth functions $\tilde{g}$, suitable for locally Lipschitz environments, facilitating the optimization process via properties like continuity, differentiability, and gradient properties, ensuring $\tilde{g}(\mathbf{x},\mu)$ converges to $g(\mathbf{x})$ as $\mu\to 0^{+}$ [22].
Our smoothing ADMM algorithm employs these properties to approximate the $\|\mathbf{z}\|_{1}$ norm effectively by using a sum of smooth functions for each $|z_{i}|$ presented in [22] as $h\mathopen{}\left(\mathbf{z},\mu\right)\mathclose{}=\sum_{i=1}^{n}f\mathopen{}\left(z_{i},\mu\right)\mathclose{}$, where:
@@ III Decentralized Smoothing ADMM
¯${\bar{\mathcal{L}}}_{\sigma_{\Psi},\sigma_{\xi},\mu}\mathopen{}\left(\mathbf{W},\mathbf{Z},\mathbf{G},\boldsymbol{\Psi},\boldsymbol{\xi}\right)\mathclose{}=\sum_{l=1}^{L}\Bigg{(}\frac{1}{2}\mathopen{}h\left(\mathbf{z}_{l},\mu\right)+\tau\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}-\\ \frac{1}{2}\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}+M_{l}P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}_{l}\right)\mathclose{}+\frac{\sigma_{\Psi}}{2}\mathopen{}\left\|\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}-\mathbf{y}^{(l)}+\frac{\boldsymbol{\Psi}_{l}}{\sigma_{\Psi}}\right\|_{2}^{2}\mathclose{}+\\ \sum_{j\in\mathcal{N}_{l},j>l}\bigg{(}\omega\hskip 1.42262pth\left(\mathbf{g}_{lj}-\mathbf{g}_{jl},\mu\right)+\frac{\sigma_{\xi}}{2}\|\mathbf{w}_{i}-\mathbf{g}_{lj}\|^{2}_{2}+\frac{\sigma_{\xi}}{2}\|\mathbf{w}_{j}-\mathbf{g}_{jl}\|^{2}_{2}\\ +\boldsymbol{\xi}_{lj}^{\text{T}}(\mathbf{w}_{l}-\mathbf{g}_{lj})+\boldsymbol{\xi}_{jl}^{\text{T}}(\mathbf{w}_{j}-\mathbf{g}_{jl})\bigg{)}\Bigg{)}$ (7)
@@
To iteratively adjust the approximation, we update the parameters $\mu$, $\sigma_{\Psi}$, $\sigma_{\xi}$ in each iteration using the following rules, where $c>0$, $d>0$ and $\beta>0$:
@@
The update process for each $\mathbf{w}_{l}$ is divided into $P+1$ sequential steps, with the $p$th element updated in its corresponding step. After simplifications, the update formula, at each agent $l$, for the $p$th for $p\leq P$ is as follows:
$w_{p,l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\textbf{Prox}_{g_{\lambda,\gamma}}\mathopen{}\left(\frac{a_{p,l}}{\Upsilon^{(l)}_{p}};\frac{M_{l}}{\Upsilon^{(l)}_{p}}\right)\mathclose{}$ (9)
@@
$a_{p,l}=-\left(\mathbf{x}^{(l)}_{:,p}\right)^{\text{T}}\left(\mathbf{X}^{(l)}_{:,<p}\mathbf{w}^{\mathopen{}\left(k+1\right)\mathclose{}}_{<p,l}+\mathbf{X}^{(l)}_{:,>p}\mathbf{w}^{\mathopen{}\left(k\right)\mathclose{}}_{>p,l}\right)\\ -\left(\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)^{\text{T}}\mathbf{X}^{(l)}_{:,p}+\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\left(\mathbf{y}^{(l)}-\mathbf{z}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)^{\text{T}}\mathbf{X}^{(l)}_{:,p}+\\ \sum_{j\in\mathcal{N}_{l},j<l}(\sigma_{\xi}^{(k+1)}g_{lj,p}^{(k)}-\xi_{lj,p}^{(k)})+\sum_{j\in\mathcal{N}_{l},j>l}(\sigma_{\xi}^{(k+1)}g_{jl,p}^{(k)}-\xi_{jl,p}^{(k)})$ (10)
@@
$\Upsilon^{(l)}_{p}=\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\mathopen{}\left\|\mathbf{X}^{(l)}_{:,p}\right\|_{2}^{2}\mathclose{}+\kappa_{\xi}^{(k+1)}$ (11)
Both MCP and SCAD admit closed-form solutions of the proximal operator [23]. For $(P+1)$th element as the penalty function is zero we have:
$w_{P+1,l}=\frac{a_{P+1,l}}{\Upsilon^{(l)}_{P+1}}$ (12)
@@
$\mathbf{z}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\operatorname*{arg\,min}_{\mathbf{z}_{l}}\sum_{j=1}^{n}\frac{1}{2}f\mathopen{}\left(z_{l,j},\mu^{\mathopen{}\left(k+1\right)\mathclose{}}\right)\mathclose{}+\mathopen{}\left(\tau-\frac{1}{2}\right)\mathclose{}\mathbf{1}_{M_{l}}^{\text{T}}\mathbf{z}_{l}$
$+\mathopen{}\left(\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}\right)\mathclose{}^{\text{T}}\mathbf{z}_{l}+\frac{\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}}{2}\mathopen{}\left\|\mathbf{z}_{l}+\mathbf{X}^{(l)}\mathbf{w}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}-\mathbf{y}^{(l)}\right\|_{2}^{2}\mathclose{}$ (13)
It can be shown that the update step of each $\mathbf{z}_{l}$ in ADMM has a closed-form solution. By grouping the last three terms of (III) together, a component-wise solution can be obtained as
$\mathbf{z}_{i}^{\mathopen{}\left(k+1\right)\mathclose{}}=\textbf{Prox}_{f\mathopen{}\left(\cdot,\mu^{\mathopen{}\left(k+1\right)\mathclose{}}\right)\mathclose{}}\mathopen{}\left(\alpha_{l,i};\frac{\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}}{2}\right)_{i=1}^{M_{l}}\mathclose{}$ (14)
@@
Updates for both $\mathbf{g}_{lj}$ and $\mathbf{g}_{jl}$ for $j\in\mathcal{N}_{l}$, $j<l$ can be performed in parallel as:
$\begin{bmatrix}\mathbf{g}_{lj}^{(k+1)}\\ \mathbf{g}_{jl}^{(k+1)}\end{bmatrix}=\operatorname*{arg\,min}_{\mathbf{g}_{lj},\mathbf{g}_{jl}}\omega\sum_{p=1}^{P+1}f({g}_{lj,p}-{g}_{jl,p},\mu^{\mathopen{}\left(k+1\right)\mathclose{}})+\frac{\sigma_{\xi}}{2}\times\\ \phantom{=}\Bigg{(}\left\|\mathbf{g}_{lj}-\mathbf{w}_{l}^{(k+1)}+\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}^{(k+1)}}\right\|^{2}_{2}+\left\|\mathbf{g}_{jl}-\mathbf{w}_{l}^{(k+1)}+\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}}\right\|^{2}_{2}\Bigg{)}$ (16)
@@
$\begin{bmatrix}\mathbf{g}_{lj}^{(k+1)}\\ \mathbf{g}_{jl}^{(k+1)}\end{bmatrix}=\\ \phantom{=}\frac{1}{2}\begin{bmatrix}\mathbf{w}_{l}^{(k+1)}-\frac{\boldsymbol{\xi}_{lj}^{(k)}}{\sigma_{\xi}^{(k+1)}}+\mathbf{w}_{j}^{(k+1)}-\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\sigma_{\xi}^{(k+1)}}\\ \mathbf{w}_{i}^{(k+1)}-\frac{\boldsymbol{\xi}_{lj}^{(k)}}{\sigma_{\xi}^{(k+1)}}+\mathbf{w}_{j}^{(k+1)}-\frac{\boldsymbol{\xi}_{jl}^{(k)}}{\kappa_{\xi}^{(k+1)}}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-\mathbf{e}\\ \mathbf{e}\end{bmatrix}$ (17)
@@
$\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}=\boldsymbol{\Psi}_{l}^{\mathopen{}\left(k\right)\mathclose{}}+\sigma_{\Psi}^{\mathopen{}\left(k+1\right)\mathclose{}}\mathopen{}\left(\mathbf{z}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}+\mathbf{X}^{(l)}\mathbf{w}_{l}^{\mathopen{}\left(k+1\right)\mathclose{}}-\mathbf{y}^{(l)}\right)\mathclose{}$ (18)
$\boldsymbol{\xi}^{(k+1)}_{lj}=\boldsymbol{\xi}^{(k)}_{lj}+\sigma^{(k+1)}_{\xi}\left(\mathbf{w}^{(k+1)}_{l}-\mathbf{g}^{(k+1)}_{lj}\right)$ (19)
$\boldsymbol{\xi}^{(k+1)}_{jl}=\boldsymbol{\xi}^{(k)}_{jl}+\sigma_{\xi}^{(k+1)}\left(\mathbf{w}^{(k+1)}_{j}-\mathbf{g}^{(k+1)}_{jl}\right)$ (20)
The proposed ADMM-based method for solving the decentralized sparse-penalized quantile regression is summarized in Algorithm 1 and the convergence is established by the following theorem.
@@
$\mathbf{X}^{\textbf{T}}\boldsymbol{\Psi}^{*}\in n\partial P_{\lambda,\gamma}\mathopen{}\left(\mathbf{w}^{*}\right)\mathclose{}$ (21a)
$\boldsymbol{\Psi}^{*}\in\partial\sum_{i=1}^{n}\rho_{\tau}\mathopen{}\left(z_{i}^{*}\right)\mathclose{}$ (21b)
$\mathbf{z}^{*}+\mathbf{X}\mathbf{w}^{*}-\mathbf{y}=0$ (21c)
@@
¯${\bar{\mathcal{L}}}_{\sigma_{\Psi}^{(K^{\prime})},\sigma_{\xi}^{(K^{\prime})},\mu^{(K^{\prime})}}\mathopen{}\left(\mathbf{W}^{(K^{\prime})},\mathbf{Z}^{(K^{\prime})},\mathbf{G}^{(K^{\prime})},\boldsymbol{\Psi}^{(K^{\prime})},\boldsymbol{\xi}^{(K^{\prime})}\right)\mathclose{}-\\ {\bar{\mathcal{L}}}_{\sigma_{\Psi}^{(K)},\sigma_{\xi}^{(K)},\mu^{(K)}}\mathopen{}\left(\mathbf{W}^{(K)},\mathbf{Z}^{(K)},\mathbf{G}^{(K)},\boldsymbol{\Psi}^{(K)},\boldsymbol{\xi}^{(K)}\right)\mathclose{}\leq\\ -S_{K^{\prime}}-S^{\prime}_{K^{\prime}}+D$ (22)
@@
$S_{K^{\prime}}=\sum_{k=K}^{K^{\prime}-1}\sum_{l=1}^{L}\eta^{(k+1)}\|\mathbf{z}_{l}^{(k+1)}-\mathbf{z}_{l}^{(k)}\|_{2}^{2}\\ +\varkappa^{(k+1)}\|\mathbf{w}_{l}^{(k+1)}-\mathbf{w}_{l}^{(k)}\|_{2}^{2}$ (23)
@@
$S_{K^{\prime}}=\sum_{k=K}^{K^{\prime}-1}\sum_{l=1}^{L}\sum_{j\in\mathcal{N}_{l},j>l}\iota^{(k+1)}\|\mathbf{q}_{lj}^{(k+1)}-\mathbf{q}_{lj}^{(k)}\|^{2}_{2}$ (24)